40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Mon Jun 11 09:26:37 2018, Disha Kapasi, Update, General, Current Driver v2 and Frequency noise of the beat note Current_Driver_Circuit_-_v2.pdfnewbeatnote.pdf
    Reply  Tue Aug 14 23:21:01 2018, johannes, Update, General, Current Driver v2 20180814_160659.jpgn_channel_driver.PNGp_channel_driver.PNGvis_mods.pdf
       Reply  Wed Aug 15 23:34:38 2018, johannes, Update, General, Cooling down 
          Reply  Mon Aug 20 02:35:06 2018, johannes, Update, General, We're cool 
Message ID: 2134     Entry time: Wed Aug 15 23:34:38 2018     In reply to: 2131     Reply to this: 2139
Author: johannes 
Type: Update 
Category: General 
Subject: Cooling down 

I started cooling down the cryostat this morning. In preparation I had flooded the inner volume with helium yesterday and let it sit overnight while pumping on the outer volume to check if the seals on the inenr viewports are still intact after several cold cycles (these are viton seals, after all). When checking the next morning, the pressure had increased by ~50% (since the volume wasn't pumped), so it looks like the viewport seals are still good. I pumped on the inner volume briefly to bring it down to ~1e-5 torr (starting pressure was 4e-6 torr) and re-flooded with only helium to ~1e-2 torr. Liquid LN2 was added at ~10am. After ten hours the cavity can temperature was passing 180 K, and I expect them to have passed 123 K in the morning, at which time I will turn on the heaters with target temps of 120 K to leave some room for the heating from the average visible laser power.

ELOG V3.1.3-